if ‘n’ is an odd integer then show that n2 – 1 is divisible than 8 ?

Grade:10

1 Answers

Nicho priyatham
8 years ago
any odd integer n=2k+1(say)k= integer
now n2-1=(n-1)(n+1)
=2k(2k+2)=4k(k+1) 
now product of any 2 concutive numbers is divisible by  2
so4k(k+1) is divisible by 8 
PLZ APPROVE IF useful and incase of any doubts ask in ans box